OpenStudy (anonymous):

Use the equation and type the ordered-pairs. Please Help y = 2x {(-1,___a0), (0,___a1), (1,___a2), (2,___a3), (3,___a4), (4,___a5)}

OpenStudy (anonymous):

y = log_2 (x) =====> when x = 1/2 y = log_2 (1/2) y = -1 ======> (1/2 , -1) y = log_2 (x) =====> when x = 1 y = log_2 (1) y = 0 ======> (1 , 0) y = log_2 (x) =====> when x = 2 y = log_2 (2) y = 1 ======> (2 , 1) y = log_2 (x) =====> when x = 4 y = log_2 (4) y = 2 ======> (4 , 2) y = log_2 (x) =====> when x = 8 y = log_2 (8) y = 3 ======> (8 , 3) y = log_2 (x) =====> when x = 16 y = log_2 (16) y = 4 ======> (16 , 4) ========

OpenStudy (anonymous):

We should be freinds on FB

OpenStudy (anonymous):

Haha ^ This guy